OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-5, 1) with a slope of 2/3 I WILL FOLLOW AND GIVE MEDAL

OpenStudy (anonymous):

y= 2/3x+13/3

OpenStudy (anonymous):

That is none of the answers

OpenStudy (anonymous):

@lacrosseplayer22

OpenStudy (anonymous):

@magbak

OpenStudy (magbak):

yes

OpenStudy (anonymous):

Help. Me

OpenStudy (anonymous):

what are the choices..??give here...

OpenStudy (anonymous):

use this eq.. (y-y1)=m(x-x1) so your eq will be... (y-1)=3/2(x+5) =>2(y-1)=3(x+5) =>2y-2=3x+15 =>3x-2y=-17 is it a choice/..???